Determine the root mean square velocity of nitrogen at a

temperature of 300 K.


1
Expert's answer
2021-07-21T14:39:50-0400

Gives

T=300k

Nitrogen

"M_N=0.028kg"

R=8.314"Jmol^-1k^-1"


"V_{RMS}=\sqrt\frac{3RT}{M_N}=\sqrt\frac{3\times8.314\times300}{0.028}=517m/sec"